# Improvise, Adapt, Overcome — Telescopic Adapters for Efficient Fine-tuning of Vision Language Models in Medical Imaging

### v1

@@ 4.5 Comparison to State-of-the-Art Methods
CLIPSeg [clipseg] (150M parameters) serves as our baseline VLSM, extending CLIP’s vision-language capabilities to segmentation tasks. Its zero-shot performance, however, is substantially low (e.g., 53.42% DSC on Kvasir-SEG), establishing a baseline that highlights the need for adaptation. While E2E fine-tuned CLIPSeg achieves competitive performance across datasets (87.69% DSC on Kvasir-SEG, 91.88% DSC on ISIC-16), it requires updating the entire parameter space. CRIS [cris] (147M parameters) employs a CNN-based CLIP encoder for region retrieval-based segmentation, achieving strong results particularly on BKAI (92.62% DSC) and ClinicDB (93.63% DSC). However, CRIS’s convolutional encoder is incompatible with adapter insertion, necessitating E2E fine-tuning and limiting its efficiency for resource-constrained deployment.
Parameter-efficient approaches offer substantial computational savings while maintaining competitive performance. I-MedSAM [imedsam] (1.6M parameters) combines frequency and LoRA adapters within the SAM framework. Despite its efficiency, I-MedSAM shows mixed results across datasets, achieving 88.05% DSC on BKAI but only 61.52% on BUSI. SAN [san] (8.4M parameters) attaches a parallel ViT network to frozen CLIP encoders, modeling segmentation as region recognition with mask proposal and attention bias branches. However, SAN demonstrates poor performance, with particularly low scores on BKAI (66.26% DSC) and BUSI (45.61% DSC), suggesting limited effectiveness for medical domain adaptation.
Existing adapter methods for VLSMs include CLIPSeg Dense Adapter (DA-VLC, 3.2M) and Shallow Adapter (SA-VLC, 4.2M) variants [vlsmadap], which apply uniform adaptation across vision, text, and conditioning pathways without depth-aware scaling. CLIPSeg DA-VLC achieves consistent performance across datasets (89.06% DSC on Kvasir-SEG, 92.02% DSC on ISIC-16), while SA-VLC shows slightly lower performance (86.80% DSC on Kvasir-SEG, 91.55% DSC on ISIC-16). Both methods require significantly more parameters than our approach while achieving inferior results. As shown in Fig. 3, our Telescopic Adapters produce segmentation masks that closely align with ground truth across diverse medical imaging modalities, demonstrating superior qualitative performance compared to existing methods.

### v2

@@ 4.5 Comparison to State-of-the-Art Methods
CLIPSeg [32] (150M parameters) serves as our baseline VLSM, extending CLIP’s vision-language capabilities to segmentation tasks. Its zero-shot performance, however, is substantially low (e.g., 53.42% DSC on Kvasir-SEG), establishing a baseline that highlights the need for adaptation. While E2E fine-tuned CLIPSeg achieves competitive performance across datasets (87.69% DSC on Kvasir-SEG, 91.88% DSC on ISIC-16), it requires updating the entire parameter space. CRIS [42] (147M parameters) employs a CNN-based CLIP encoder for region retrieval-based segmentation, achieving strong results particularly on BKAI (92.62% DSC) and ClinicDB (93.63% DSC). However, CRIS’s convolutional encoder is incompatible with adapter insertion, necessitating E2E fine-tuning and limiting its efficiency for resource-constrained deployment.
Parameter-efficient approaches offer substantial computational savings while maintaining competitive performance. I-MedSAM [45] (1.6M parameters) combines frequency and LoRA adapters within the SAM framework. Despite its efficiency, I-MedSAM shows mixed results across datasets, achieving 88.05% DSC on BKAI but only 61.52% on BUSI. SAN [46] (8.4M parameters) attaches a parallel ViT network to frozen CLIP encoders, modeling segmentation as region recognition with mask proposal and attention bias branches. However, SAN demonstrates poor performance, with particularly low scores on BKAI (66.26% DSC) and BUSI (45.61% DSC), suggesting limited effectiveness for medical domain adaptation.
Existing adapter methods for VLSMs include CLIPSeg Dense Adapter (DA-VLC, 3.2M) and Shallow Adapter (SA-VLC, 4.2M) variants [6], which apply uniform adaptation across vision, text, and conditioning pathways without depth-aware scaling. CLIPSeg DA-VLC achieves consistent performance across datasets (89.06% DSC on Kvasir-SEG, 92.02% DSC on ISIC-16), while SA-VLC shows slightly lower performance (86.80% DSC on Kvasir-SEG, 91.55% DSC on ISIC-16). Both methods require significantly more parameters than our approach while achieving inferior results. As shown in Fig. 3, our Telescopic Adapters produce segmentation masks that closely align with ground truth across diverse medical imaging modalities, demonstrating superior qualitative performance compared to existing methods.